# Bridging Relevance and Reasoning: Rationale Distillation in Retrieval-Augmented Generation

### v1

@@ 3.6 Transferability Analysis across Generators
We conduct experiments on two datasets of different tasks, NQ and MMLU, using three different generators (Llama3.1-8b-instruct Touvron et al. (2023), qwen2.5-14b-instruct Yang et al. (2024), and gpt4o-mini Achiam et al. (2023)) to validate the transferability of our method, as shown in Table 3. We can find: (1) RADIO maintains its effectiveness across different generators, consistently enhancing the performance of the original RAG pipeline. This demonstrates RADIO’s strong transferability with various generators. (2) Comparing different generators reveals that RADIO’s performance gains are more pronounced with smaller, less capable generators. Specifically, when the generators are Llama3.1, Qwen2.5, and GPT4o-mini, RADIO achieves EM improvements of 8.72%, 6.28%, and 3.74%, respectively, and F1 improvements of 6.82%, 6.00%, and 4.22%. This is because as the generator’s capability increases and approaches the upper performance limits of the RAG pipeline, further enhancing the pipeline becomes increasingly challenging, resulting in a smaller improvement.
@@ 3.7 Ablation Study
To explore the specific impact of rationale and retrieval score, we design the following variants: (1) w/o ALL: Base reranker without fine-tuning. Do not introduce rationale or retrieval score. (2) w/o Retrieval: Ranking documents and fine-tuning reranker only based on the rationale scores. (3) RADIO: Fine-tuning reranker based on both rationale and retrieval scores.
@@ 3.8 Hyperparameter Analysis
Figure 2 visualizes the performance of RADIO across different integration coefficients $\alpha$ on NQ and MMLU. The x-axis represents the integration coefficients $\alpha$ and the y-axis represents the evaluation metrics EM (red) or F1 (blue). For the MMLU dataset, we present results for two representative categories: Humanities and STEM. The trends for other categories align with those observed in Humanities. Complete experimental results are provided in Appendix A.3 for reference. From the figure, we can draw the following conclusions: (1) As $\alpha$ increases, the metrics on the NQ dataset and most categories of the MMLU dataset exhibit a trend of first rising and then falling, with the optimal range for $\alpha$ being between 0.3 and 0.7. This demonstrates the complementary nature of rationale scores and retrieval scores, which together form an optimal signal for preference alignment. (2) When $\alpha=0$, the RAG performance is suboptimal because document scoring relies entirely on retrieval scores, focusing solely on query-document relevance while ignoring whether the document supports the generator in answering the query. Conversely, when $\alpha=1$, the performance is still not optimal, as it completely disregards retrieval relevance, leading to a mismatch between the fine-tuning dataset and the training data, which negatively affects model performance. (3) The trends for STEM differ from those of other MMLU categories, showing an opposite pattern. This is likely due to the significant distributional differences between STEM and Humanities/Social Sciences, resulting in a "seesaw effect" as observed in the figure. This phenomenon is also reflected in REPLUG Shi et al. (2023), where the improvement in the STEM category is weaker compared to other categories.
3.9 Case Study
Figure 3: Case study on NQ dataset.
To intuitively illustrate the effectiveness of RADIO, we select examples from the NQ dataset to compare the documents reranked by RADIO with those reranked by a novel baseline, DPA-RAG, as well as their responses. In Figure 3 Example 1, the query asks, "Which state is the richest state in Nigeria?" RADIO successfully ranks information about Lagos State’s economic and financial status, which relates to the correct answer, among the top-3 documents. In contrast, DPA-RAG fails to identify documents relevant to answering the query, and cannot provide a correct response. In Figure 3 Example 2, the query is, "Who is the highest-selling R&B artist of all time?" RADIO prioritizes documents containing information about the correct answer, Michael Jackson, and effectively highlights key terms such as "R&B" and "best-selling." However, DPA-RAG misinterprets the query’s constraints, retrieving documents that either overlook the R&B artist specification or fail to consider the time span, resulting in an incorrect response. These examples demonstrate that RADIO enhances the RAG pipeline by providing more efficient and accurate reranking. It selects contextually appropriate documents, enabling the generator to infer the correct answers effectively.
4 Related Work
4.1 Retrieval-Augmented Generation
Large language models (LLMs) have demonstrated groundbreaking performance across numerous NLP tasks but still face challenges such as hallucination and outdated knowledge Gao et al. (2023). To address these issues, retrieval-augmented generation (RAG) has been introduced Fan et al. (2024). RAG retrieves relevant information from external knowledge bases and incorporates it as contextual input to the generator (LLM), enhancing the accuracy and reliability of the generated responses. The typical RAG pipeline can be divided into several key components: query rewriter, retriever, reranker, and generator. The query rewriter Wang et al. (2023) modifies and expands the original query to improve retrieval recall, ensuring more relevant documents are retrieved. The retriever Chen et al. (2024) fetches relevant documents based on the query. Dense retrievers generally outperform sparse retrievers in this step. To integrate contextual information more effectively and identify documents more relevant to the query, rerankers Moreira et al. (2024a) with larger models and greater complexity are introduced to reorder the retrieved documents compared to retrievers. Finally, the generator—usually a powerful LLM such as GPT-4 Achiam et al. (2023) or Llama Touvron et al. (2023)—uses the query and the top-$k$ documents from the reranker to generate the final response.
4.2 Preference Alignment
To further improve LLMs, preference alignment is often performed after the initial pretraining phase Jiang et al. (2024). Approaches such RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2024) are proposed to align the output of LLMs more closely with human preferences. DPO transforms tasks into classification problems, achieving high computational efficiency and strong performance. In the context of RAG, several works can be transformed to address the preference alignment challenge across RAG components. REPLUG Shi et al. (2023) improves RAG pipelines involving black-box LLMs by using the probability of the LLM generating the correct answer as a signal to determine document importance. Similarly, RRR Ma et al. (2023) uses metrics based on the quality of the LLM’s generated response as a signal to evaluate a document’s utility. On the other hand, ARL2 Zhang et al. (2024) prompts LLMs to generate self-guided relevance labels for fine-tuning retriever, and DPA-RAG Dong et al. (2024) introduces a bidirectional alignment strategy to mitigate preference inconsistencies in RAG pipelines. In this work, we focus on optimizing the reranker within RAG. Our goal is to enable the reranker to effectively identify supportive documents well-suited for the generator, whether black-box or open-source, facilitating the production of accurate outputs. In addition, BGM Ke et al. (2024) trains a bridge model between retriever and LLMs to transform the retrieved information into the format LLMs prefer, which can be used with RADIO in combination.
@@ 5 Conclusion
In this paper, we propose a novel and practical preference alignment framework, RADIO, with rationale distillation in retrieval-augmented generation. First, we introduce a rationale extraction method to extract the rationales necessary for answering queries with LLMs. Next, a rationale-based alignment is proposed to rerank documents based on extracted rationales and fine-tune rerankers. Extensive experiments are conducted to verify the effectiveness of our proposed method compared to other state-of-the-art baselines.
@@ Appendix A Appendix
Figure 4: More results of hyperparameter analysis on MMLU.
@@ A.2 Baselines
  • ARL2 Zhang et al. (2024): ARL2 introduces a method to use LLMs as supervisor to generate self-guided relevance labels for fine-tuning retriever.
@@
A.3 More Results on Hyperparameter Analysis
Table 5: Prompts for Open-domain datasets.
@@ A.3 More Results on Hyperparameter Analysis
A.4 Prompts
In this section, we detail the prompts we used in the experiments. For Open-domain QA datasets NQ and TriviaQA, the prompts are as shown in Table 5. For Multi-choice dataset MMLU, the prompts are shown in Table 6.

### v2

@@ 4.6.1 Multi-choice Questions
4.6.2 Multi-hop QA
Table 4: Experimental results on Musique. EM and F1 are reported as the metrics. The source datasets used to fine-tune rerankers are Open-domain QA datasets NQ and TriviaQA. $\Delta$ represents the relative improvement of RADIO over Base method.
Method NQ (Source Dataset) TriviaQA (Source Dataset)
EM F1 EM F1
Base 0.0662 0.1297 0.0662 0.1297
Altlas 0.0658 0.1336 0.0541 0.1184
REPLUG 0.0608 0.1269 0.0559 0.1188
RRR 0.0658 0.1315 0.0592 0.1207
ARL2 0.0641 0.1309 0.0674 0.1330
DPA-RAG 0.0629 0.1307 0.0690 0.1337
RADIO (Ours) 0.0699 0.1371 0.0695 0.1347
$\Delta$uparrow5.59$\uparrow$5.71$\uparrow$4.98$\uparrow$3.86%
We also conduct experiments on the multi-hop dataset Musique with reranker bge-reranker-base, and the experimental results are shown in Table 4. Based on the experimental results, we can draw the following conclusions:
  • First, RADIO achieves significant performance improvements in all evaluation metrics (EM and F1), demonstrating its task generalization ability in multi-hop reasoning tasks. On the NQ source dataset, RADIO achieves an EM value of 0.0699 and an F1 value of 0.1371, with improvements of 5.59% and 5.71% over the Base method, respectively. On the TriviaQA source dataset, RADIO’s EM value is 0.0695, and F1 value is 0.1347, showing improvements of 4.98% and 3.86% over the Base method.
  • Secondly, compared to other baseline methods, RADIO demonstrates a more balanced and consistent advantage across both the NQ and TriviaQA source datasets. Notably, when using NQ as the source dataset, only RADIO shows a positive improvement in the EM metric, while all other baseline methods experience varying degrees of negative degradation.
@@ 4.7 Transferability Analysis across Generators
We conduct experiments on two datasets of different tasks, NQ and MMLU, using three different generators (Llama3.1-8b-instruct Touvron et al. (2023), qwen2.5-14b-instruct Yang et al. (2024), and gpt4o-mini Achiam et al. (2023)) with bge-reranker-base fine-tuned on NQ to validate the transferability of our method, as shown in Table 3. We can find that (1) RADIO maintains its effectiveness across different generators, consistently enhancing the performance of the original RAG pipeline. This demonstrates RADIO’s strong transferability with various generators. (2) Comparing different generators reveals that RADIO’s performance gains are more pronounced with smaller, less capable generators. Specifically, when the generators are Llama3.1, Qwen2.5, and GPT4o-mini, RADIO achieves EM improvements of 8.72%, 6.28%, and 3.74%, respectively, and F1 improvements of 6.82%, 6.00%, and 4.22%. This is because as the generator’s capability increases and approaches the upper performance limits of the RAG pipeline, further enhancing the pipeline becomes increasingly challenging, resulting in a smaller improvement.
@@ 4.8 Ablation Study
To explore the specific impact of rationale and retrieval score, we design the following variants with reranker bge-reranker-base:
  • w/o ALL: Base reranker without fine-tuning. Do not introduce rationale or retrieval score.
  • w/o Retrieval: Ranking documents and fine-tuning reranker only based on the rationale scores.
  • RADIO: Fine-tuning reranker based on both rationale and retrieval scores.
@@ 4.9 Hyperparameter Analysis
Figure 2 visualizes the performance of RADIO across different integration coefficients $\alpha$ on NQ and MMLU. The x-axis represents the integration coefficients $\alpha$ and the y-axis represents the evaluation metrics EM (red) or F1 (blue). For the MMLU dataset, we present results for two representative categories: Humanities and STEM. The trends for other categories align with those observed in Humanities. Complete experimental results are provided in Appendix A.9 for reference. From the figure, we can draw the following conclusions: (1) As $\alpha$ increases, the metrics on the NQ dataset and most categories of the MMLU dataset exhibit a trend of first rising and then falling, with the optimal range for $\alpha$ being between 0.3 and 0.7. This demonstrates the complementary nature of rationale scores and retrieval scores, which together form an optimal signal for preference alignment. (2) When $\alpha=0$, the RAG performance is suboptimal because document scoring relies entirely on retrieval scores, focusing solely on query-document relevance while ignoring whether the document supports the generator in answering the query. Conversely, when $\alpha=1$, the performance is still not optimal, as it completely disregards retrieval relevance, leading to a mismatch between the fine-tuning dataset and the training data, which negatively affects model performance. (3) The trends for STEM differ from those of other MMLU categories, showing an opposite pattern. This is likely due to the significant distributional differences between STEM and Humanities/Social Sciences, resulting in a "seesaw effect" as observed in the figure. It is worth noting that this phenomenon is also reflected in REPLUG Shi et al. (2023), where the improvement in the STEM category is weaker compared to other categories.
@@ 5 Conclusion
In this paper, we propose a novel and practical preference alignment framework, RADIO, with rationale distillation in retrieval-augmented generation. First, we introduce a rationale extraction method to extract the rationales necessary for answering queries with LLMs. Next, a rationale-based alignment is proposed to rerank documents based on extracted rationales and fine-tune rerankers. Extensive experiments on three tasks across four datasets are conducted to validate the effectiveness of our proposed method against state-of-the-art baselines and demonstrate its strong transferability. We also release our code online to facilitate the following research and ease of reproduction.
@@ 6 Limitations
Acknowledgments
This research was partially supported by Research Impact Fund (No.R1015-23), Collaborative Research Fund (No.C1043-24GF), and Huawei (Huawei Innovation Research Program, Huawei Fellowship).
@@ A.1 Dataset Desctiptions
  • Musique: MuSiQue is a multi-hop question answering dataset created using a bottom-up approach to ensure connected reasoning, with 25,000 2–4 hop questions.
@@ A.2 Baselines
  • ARL2 Zhang et al. (2024): ARL2 introduces a method to use LLMs as supervisor to generate self-guided relevance labels (e.g., “Not relevant”, “relevant”) for fine-tuning retriever.
@@
A.3 Detailed Introduction to Backbone Rerankers
To validate the generality and adaptability of RADIO, we select three different rerankers as the backbone models for our experiments: (1) gte-base Li et al. (2023b): a reranker model proposed by Alibaba DAMO Academy, with 109M parameters and 768 embedding dimensions. (2) gte-large Li et al. (2023b): the larger version of gte-base, with 335M parameters and 1024 embedding dimensions. (3) bge-reranker-base Xiao et al. (2023): a powerful cross-encoder architecture reranker proposed by Beijing Academy of Artificial Intelligence, with 278M parameters.
A.4 Prompts
Table 6: Prompts for QA datasets.
@@ A.4 Prompts
In this section, we detail the prompts we used in the experiments. For Open-domain QA datasets (NQ and TriviaQA) and Multi-hop QA dataset (Musique), the prompts are as shown in Table 6. For Multi-choice dataset MMLU, the prompts are shown in Table 7.
A.5 Sample Efficiency Analysis of Fine-tuning
To understand the impact of fine-tuning samples on RAG performance, we conduct additional experiments to evaluate RADIO’s performance under a few-shot learning setup using varying amounts of fine-tuning samples (0, 4,000, 8,000, 12,000, 16,000, and 20,000 instances) on the NQ dataset with bge-reranker-base. The results are shown in Table 8.
Table 8: Experimental results on NQ with different numbers of fine-tuning samples.
Sample Num. EM F1
0 0.3371 0.4603
4,000 0.3526 0.4843
8,000 0.3634 0.4896
12,000 0.3582 0.4842
16,000 0.3604 0.4889
20,000 0.3665 0.4917
We observe that even with a relatively small number of fine-tuning samples (e.g., 4,000 or 8,000), RADIO yields substantial improvements over the zero-shot baseline. Notably, with 8,000 samples the EM score increases by 7.8% and the F1 score by 6.4%, clearly demonstrating the model’s capacity to benefit from limited task-specific data. Moreover, as the number of fine-tuning instances increases, the overall performance tends to improve, suggesting that additional annotated data can further enhance the retrieval and generation capabilities of the system. Although there are minor fluctuations—such as a slight dip in performance at 12,000 and 16,000 samples—these variations may stem from differences in data distribution or optimization dynamics.
A.6 More Analysis on Training Hyperparameters
To understand the specific impact of two training hyperparameters (i.e., the shift $n$ in Top-k shifted by N and the number of negative samples $m$) on performance, we conduct the following experiments on NQ based on bge-reranker-base:
  • Varying $m$: We fix $n=3$ and vary $m$ over the values 2,4,6, and 8.
  • Varying $n$: We fix $m=6$ and vary $n$ over the values 1,2,3,4, and 5.
The detailed experimental results are presented in Table 9 and Table 10, respectively. From the results, we can find the following points: (1) Impact of Negative Samples $m$: Variations in the number of negative samples $m$ have a relatively modest effect on RADIO’s performance. The results suggest that the model is robust to changes in $m$, as only marginal performance differences are observed when $m$ is varied from 2 to 8. (2) Impact of the shift parameter $n$: In contrast, the shift parameter $n$ in the Top-k shifted by N mechanism plays a more critical role. Our experiments show that smaller values of $n$ yield better performance, while larger $n$ values lead to a notable drop in both EM and F1 scores. We hypothesize that as $n$ increases, the task of mining hard negatives becomes more challenging, which in turn hampers the fine-tuning process of RADIO.
Table 9: Hyperparameter analysis on the number of negative samples $m$.
Shift ($n$) Negative Samples $m$ EM F1
3 2 0.3615 0.4886
3 4 0.3601 0.4888
3 6 0.3665 0.4917
3 8 0.3604 0.4895
Figure 3: Case study on NQ dataset.
Table 10: Hyperparameter analysis on the shift $n$ in Top-k shifted by N
Shift ($n$) Negative Samples $m$ EM F1
1 6 0.3654 0.4918
2 6 0.3554 0.4826
3 6 0.3665 0.4917
4 6 0.3579 0.4868
5 6 0.3587 0.4887
A.7 Case Study
To intuitively illustrate the effectiveness of RADIO, we select examples from the NQ dataset to compare the documents reranked by RADIO with those reranked by a novel baseline, DPA-RAG, as well as their responses. In Figure 3 Example 1, the query asks, "Which state is the richest state in Nigeria?" RADIO successfully ranks information about Lagos State’s economic and financial status, which relates to the correct answer, among the top-3 documents. In contrast, DPA-RAG fails to identify documents relevant to answering the query, and cannot provide a valid response. In Figure 3 Example 2, the query is, "Who is the highest-selling R&B artist of all time?" RADIO prioritizes documents containing information about the correct answer, Michael Jackson, and effectively highlights key terms such as "R&B" and "best-selling." However, DPA-RAG misinterprets the query’s constraints, retrieving documents that either overlook the R&B artist specification or fail to consider the time span, resulting in an incorrect response. These examples demonstrate that RADIO enhances RAG by providing a more efficient and accurate reranking. It selects contextually appropriate documents, enabling the generator to infer correct answers.
A.8 More Results on MMLU
In this section, we give the complete experimental results on MMLU. Specifically, we use the NQ dataset and TriviaQA dataset as source dataset to fine-tune rerankers and evaluate them in MMLU. The results are shown in Table 11.
Method MMLU (Source Dataset NQ) MMLU (Source Dataset TriviaQA)
Humanities Social STEM Other ALL Humanities Social STEM Other ALL
RADIO (Ours) 0.4172 0.7013 0.5080 0.6717 0.5562 0.4230 0.6942 0.5090 0.6678 0.5559
A.9 More Results on Hyperparameter Analysis
Figure 4: More results of hyperparameter analysis on MMLU.